# NSmark: Null Space Based Black-box Watermarking Defense Framework for Pre-trained Language Models

### v1

@@ 3.4.3 Watermark Verification
To better defend against possible attacks, NSmark  uses two metrics together to verify the ownership: Watermark Extracting Rate (WER) and NSMD. Model owner needs to submit $key=(sig,E,N)$ to Certification Authority (CA). CA generates $t,D_{V},sm$ using $sig$. Input $D_{V}$ to the suspicious model $f_{susp}$ to get the output vector $A_{susp}$, and pass $A_{susp}$ through $E$ to get the mapped vector $O_{susp}$. Then WER is calculated after despread spectrum. Despread spectrum is the inverse process of spread spectrum (detailed process in Appendix A.5.2). At last the signature is extracted as $sig^{\prime}=\{a^{\prime}_{i}|a^{\prime}_{i}\in\{-1,0,1\},i\in[0,n-1]\}$. WER is defined as the proportion of $sig$ and $sig^{\prime}$ having the same value:
$\text{WER}=\frac{1}{n}\sum_{i=0}^{n-1}\left[a_{i}=a_{i}^{\prime}\right].$ (4)
@@
NSMD is calculated using $A_{susp}$ and $N$ by Equation 1. We define two thresholds, and whether WER$>T_{W}$ will be firstly verified. If it fails, whether NSMD$<T_{N}$ will be further considered in case of LL-LFEA.
@@ 4.1 Experimental Setup
Datasets. We use WikiText-2 Merity et al. (2017) for pre-training and watermark embedding. To evaluate the performance on downstream tasks, we select many text classification datasets: SST-2 and SST-5 Socher et al. (2013) for sentiment analysis, Lingspam Sakkis et al. (2003) for spam detection, OffensEval Zampieri et al. (2019) for offensive language identification, and AG News Zhang et al. (2015) for news classification.
@@
Table 1: Effectiveness of NSmark  on different PLMs.
Models. For PLMs, we use the base versions of BERT Kenton and Toutanova (2019), RoBERTa Liu (2019), DeBERTa He et al. (2020) and XLNet Yang (2019). All pre-trained weights are from HuggingFace [^1] https://huggingface.co/. The extractor network is a three-layer linear network with hidden layers of $2048$ and $1024$ neurons. The input dimension matches the output dimension of the PLM. The output dimension matches the size of $sig_{sm}$.
Watermark settings and training details. We select a string containing owner information as the message $m$, e.g., "BERT is proposed by Google in 2018". The length of $sig$ is $256$ and is spread-spectrum with a spreading factor $k=3$, resulting in a $768$-bit $sig_{sm}$. We use the labeled downstream dataset SST-2 as the candidate pool $D_{NS}$. $q$, the length of $D_{V}$, is $1500$. The trigger is inserted to random positions for $5$ times in trigger set. When performing watermark embedding, $\lambda_{1}=0.5$ and $\lambda_{2}=0.2$ in $L_{Extractor}$ and $L_{wm}$. Batchsize is $4$, and learning rates for both $f_{wm}$ and $E$ are $10^{-4}$. $f_{wm}$ and $E$ are trained alternately for $10$ epochs. When fine-tuning on downstream tasks, the learning rate is $2\times 10^{-5}$ and batchsize is $8$ for $3$ epochs.
Metrics. As mentioned before, We define two metrics to verify the model’s identity: WER and NSMD. Besides, we adopt accuracy (ACC, in %) to measure the performance of PLM on downstream tasks.
4.2 Performance Evaluation
@@ 4.2.1 Effectiveness
Effectiveness means that the watermark can achieve expected effect in verification. Ideally, $sig^{\prime}$ extracted from the watermarked model should be consistent with the original $sig$, and the output matrix of $f_{wm}$ for $D_{V}$ should completely match $N$ stored in the $key$, which means WER = $1$ and NSMD = 00. Table 1 shows the results of $f_{wm}$ embedded with watermark and $f_{clean}$ without watermark. It can be seen that for different watermarked PLMs, WER is $1$, and NSMD is close to 00. This proves the effectiveness of NSmark. Comparing the values of $f_{wm}$ and $f_{clean}$ shows that WER and NSMD will change obviously after the watermark is embedded. Although NSMD of different PLMs is different in value, they are all far from 00. Through these results, we can preliminarily define verification thresholds of WER and NSMD as $T_{W}=0.6$ and $T_{N}=43$, which are 0.6×0.6$\times$0.6 × the average gaps. Thresholds can be further adjusted according to different models and task types.
@@ 4.2.2 Reliability
Figure 5: The impact of the correctness of trigger $t$ and signature $sig$ on WER and NSMD.
@@
For $f_{wm}$, (1) when trigger is wrong ($t_{w}$) and signature is correct ($sig_{c}$), WER = 0, NSMD $>T_{N}$. This means that $f_{wm}$ has learned the relationship between $sig$ and $t$. Whether $t$ is correct determines whether $f_{wm}$ can produce the expected output, which in turn affects both the calculation of WER and NSMD. (2) When trigger is correct ($t_{c}$) and signature is wrong (considering the most dangerous scenario $sig_{w}$ only consists of $\{-1,1\}$), WER $\approx 0.5$. This is because $t_{c}$ leads to the right $sig^{\prime}$, and its expectation of WER with a random $\{-1,1\}$ string is $0.5$. As the output matrix is correctly generated by $f_{wm}$ based on $t_{c}$, NSMD = 00 in this case. (3) When both trigger and signature are wrong ($t_{w}$ and $sig_{w}$), WER = 00, NSMD > $T_{N}$, indicating that the watermark cannot be correctly verified without providing the correct key. (4) For model without embedded watermarks $f_{clean}$, watermarks cannot be extracted even if the correct key is provided.
@@
Table 2: WER of different extractor $E$.
Wrong extractor $E$. Since $E$ is not involved in the calculation of NSMD, we only analyze the impact of $E$ on WER. As shown in Table 2, when $E$ is wrong ($E_{w}$), WER is close to 00, indicating that wrong $E$ is unable to extract the watermark.
@@
Table 3: NSMD of different null space matrix $N$.
Metric BERT RoBERTa DeBERTa XLNet
$f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$
WER 1.00 0.27 1.00 0.07 1.00 0.27 1.00 0.00
Table 4: Impact of LL-LFEA on watermark performance.
Wrong null space $N$. Since $N$ is not involved in the calculation of WER, we only analyze its impact on NSMD. As shown in Table 3, $N_{r}$ is a randomly generated matrix with the same dimension as $N$ and each element is distributed between $[0,1]$. It can be seen that NSMD is very large at this time. However, if the attacker knows the watermark algorithm, a $N_{s}$ with extremely small elements can be generated. In this case NSMD might meet the verification requirements. This indicates that NSMD cannot be used independently to verify watermark.
4.2.3 Fidelity
@@
Table 5: The performance of models w/o watermark fine-tuned on downstream tasks.
We hope that NSmark  does not affect the performance on original tasks. Thus we add a downstream network to $f_{wm}$, and fine-tune the whole model $F_{wm}$ with downstream dataset. $F_{clean}$ without watermark is fine-tuned as baseline. Table 5 shows that watermark has almost no impact on the performance of the model on the original task.
@@ 4.2.4 Defense against LL-LFEA
Defense against LL-LFEA. When designing NSmark, we focus on resisting LL-LFEA, and propose null space verification using NSMD. Table 4 shows the impact of the LL-LFEA on the watermark verification, where $f_{LL-LFEA}$ denotes the model $f_{wm}$ attacked by LL-LFEA. Experiments show that after LL-LFEA, WER drops significantly as we discussed in §$\S$ 3.1, but NSMD is still close to 00, proving that NSMD is an effective indicator for LL-LFEA. Furthermore, after applying LL-LFEA, the attacker may add a network to $f_{LL-LFEA}$ and fine-tune it for downstream tasks. We present results in Appendix B.1.
Recovery of WER. In LFEA Li et al. (2023a), a method is proposed to recover the watermark. We revise this method to recover $f_{rec}$ from $f_{LL-LFEA}$. Specifically, assume that the output matrix of $f_{wm}$ is $A_{1(n\times m)}$ as Proof A.1. After attacked with $Q_{(n\times n)}$, the output matrix turns to $A_{2}=Q\times A_{1}$. Therefore, an estimate of $Q$ can be obtained as $Q^{\prime}=A_{2}\times A_{1}^{-1}$. If $m\neq n$, then $A_{1}$ is not reversible and $Q^{\prime}=A_{2}\times A_{1}^{T}\times(A_{1}\times A_{1}^{T})^{-1}$. Then we perform an anti-attack transformation on $f_{LL-LFEA}$, that is, multiply all the outputs of $f_{LL-LFEA}$ by $Q^{\prime}$ to get $f_{rec}$. Figure 6 shows that after recovery, WER is significantly improved, indicating that such linear attacks are recoverable. In all cases, NSMD is quite small, proving that NSMD is invariant to LL-LFEA. In the recovery algorithm in Li et al. (2023a), both the attacker and the model owner might use such an algorithm to claim that they are the owner of the model, which will cause verification ambiguity. However, our proposed NSMD is invariant under LL-LFEA, thus as long as the timestamp information is added to $key$ tuple, the ownership can be reliably verified according to the time sequence of the model and $key$ release.
Figure 6: Changes of WER and NSMD before and after LL-LFEA attack and recovery.
@@ 4.2.5 Robustness
The robustness of watermark refers to whether the watermark can be effectively verified after watermark removal attacks. Next, we will analyze the robustness of NSmark  against fine-tuning, pruning, fine-pruning and overwriting attacks.
Robustness against fine-tuning. Table 5 shows the WER and NSMD results after fine-tuning on downstream tasks. $F_{wm}$ and $F_{clean}$ are obtained same as §$\S$ 4.2.3. In most cases, WER is still very high, indicating that the embedded $sig$ can still be effectively extracted after downstream fine-tuning. However, WER of RoBERTa on Lingspam task is relatively low. This is because samples in this dataset are too long to process, and the output vectors after fine-tuning have large changes, which together lower the WER. Nevertheless, compared with $F_{clean}$, there is still obvious discrimination. Therefore, for complex tasks, the verification threshold $T_{W}$ can be slightly lowered.
Figure 7: Impact of pruning attacks on watermark performance.
Robustness against pruning and fine-pruning. Pruning is a commonly used model compression method and is often used to destroy the watermark embedded in the model. Referring to Han et al. (2015); Shao et al. (2024), we sort the parameters of each layer in PLM, then set different fractions of parameters with the smaller absolute value to 00. Figure 7 shows that when the pruning rate is less than or equal to $0.8$, WER is very high. When the pruning rate is less than or equal to $0.6$, NSMD does not change significantly, and even when the pruning rate is as high as $0.9$, NSMD is still distinguishable. This shows that the embedded watermark is robust to pruning attack.
Usually, pruning will affect the performance of the model on the original task, and the original task accuracy will be restored through fine-tuning (fine-pruning). We present corresponding results in Appendix B.2.
@@
Table 6: Impact of overwrite attacks on watermark performance.
Robustness against overwriting. Overwriting means attacker embeds his own watermark into a model that has already been watermarked in the same way. This may destroy the original watermark. We simulate this process to obtain $f_{ow}$, then add a downstream network and fine-tune to obtain $F_{ow}$. We test the original watermark as shown in Table 6. Overwriting attack has little effect on ACC and WER except on Lingspam. Meanwhile, it has a similar impact on NSMD as fine-tuning.

### v2

@@ 3.4.3 Watermark Verification
To effectively defend attacks, NSmark  uses two metrics together to verify ownership: WER and NSMD. Model owner needs to submit $key=(sig,E,N)$ to the Certification Authority (CA). CA generates $t,D_{V},sm$ using $sig$. Input $D_{V}$ to the suspicious model $f_{susp}$ to get the output vector $A_{susp}$, and pass $A_{susp}$ through $E$ to get the mapped vector $O_{susp}$. Then WER is obtained from the despread spectrum.
WER is defined from comparing bits in $sig$ and $sig^{\prime}$:
$\text{WER}=\frac{1}{n}\sum_{i=0}^{n-1}\left[a_{i}=a_{i}^{\prime}\right],$ (4)
@@
NSMD is calculated using $A_{susp}$ and $N$ by Equation 1. We define two thresholds, and whether WER$>T_{W}$ will be first verified. If it fails, whether NSMD$<T_{N}$ will be further considered in the case of LL-LFEA.
@@ 4.1 Experimental Setup
Datasets. We use WikiText-2 (Merity et al., 2017) for pre-training and watermark embedding. To evaluate the performance on downstream tasks, we select many text classification datasets: SST-2 and SST-5 (Socher et al., 2013) for sentiment analysis, Lingspam (Sakkis et al., 2003) for spam detection, OffensEval (Zampieri et al., 2019) for offensive language identification and AG News (Zhang et al., 2015) for news classification.
Table 1: Effectiveness of NSmark  on different LMs. $f_{wm}$ means watermarked model and $f_{clean}$ is not watermarked. WER=$1.00$ indicates that the signature information can be accurately extracted, and NSMD has obvious differentiation between $f_{wm}$ and $f_{clean}$.
@@
Models. For LMs, we use the base versions of BERT (Kenton & Toutanova, 2019), RoBERTa (Liu, 2019), DeBERTa (He et al., 2020) and XLNet (Yang, 2019) for main results. Llama-2-7B (Touvron et al., 2023), GPT-2 (Radford et al., 2019) and the large version of BERT and RoBERTa are also used in supplementary experiments. All pretrained weights are from HuggingFace. [^1] https://huggingface.co/ The extractor network is a three-layer linear network with hidden layers of neurons $2048$ and $1024$. The input dimension matches the output dimension of the LM. The output dimension matches the size of $sig_{sm}$.
Watermark settings and training details. We select a string containing owner information as the message $m$, for example, ”BERT is proposed by Google in 2018”. The length of $sig$ is $256$ and then spread spectrum by a factor $k=3$, resulting in a $768$-bit $sig_{sm}$. SST-2 is used as the candidate pool $D_{NS}$. $q$, the length of $D_{V}$, is $1500$. The trigger is inserted into random positions for $5$ times in the trigger set. When performing watermark embedding, $\lambda_{1}=0.5$ and $\lambda_{2}=0.2$ in $L_{Extractor}$ and $L_{wm}$. The batchsize is $4$, and the learning rates for both $f_{wm}$ and $E$ are $10^{-4}$. $f_{wm}$ and $E$ are trained alternately for the $10$ epochs. When fine-tuning downstream tasks, the learning rate is $2\times 10^{-5}$ and the batchsize is $8$ for $3$ epochs.
Metrics. As mentioned before, two metrics are defined to verify the identity of the model: WER and NSMD. Besides, we adopt accuracy (ACC, in %) to measure the performance of LM on downstream tasks.
4.2 LL-LFEA Attack Evaluation
We select the effective word embedding-based watermarking scheme (EmbMarker) (Peng et al., 2023) and NSmark  (without NSMD) as victims to study the effectiveness of LL-LFEA. The attack results on EmbMarker are shown in Appendix B.1 and the results on NSmark  (without NSMD) are shown in Table 4. Though EmbMarker can pass through the attack of RedAlarm (Zhang et al., 2023), after the LL-LFEA attack, all the metrics of EmbMarker are very close to those of the original model without watermark. At the same time, LL-LFEA has little degradation on original model performance. This fully demonstrates the effectiveness of LL-LFEA on existing watermarking schemes.
4.3 NSmark  Performance Evaluation
We analyze the main experimental results about NSmark. For more results on computational cost and other more detailed experiments, please refer to Appendix B.
@@ 4.3.1 Effectiveness
Effectiveness means that the watermark can achieve the expected effect during verification. Ideally, $sig^{\prime}$ extracted from the watermarked model should be consistent with the original $sig$, and the output matrix of $f_{wm}$ for $D_{V}$ should match completely $N$ stored in $key$, which means WER = $1$ and NSMD = 00. Table 1 shows the results of $f_{wm}$ embedded with watermark and $f_{clean}$ without watermark. It can be seen that for different watermarked LMs, WER is $1$ and NSMD is close to 00. This shows the effectiveness of NSmark. Comparison of the values of $f_{wm}$ and $f_{clean}$ shows that WER and NSMD will obviously change after the watermark is embedded. Although NSMD of different LMs has different values, they are all far from 00. Through these results, we can preliminarily define verification thresholds of WER and NSMD as $T_{W}=0.6$ and $T_{N}=43$, which are 0.6×0.6$\times$0.6 × the average gaps. Thresholds can be further adjusted according to different models and task types. We also study the watermark effectiveness on larger size of the models in Table 10, which demonstrates the scalability of NSmark.
@@ 4.3.2 Reliability
Figure 5: The impact of the correctness of trigger $t$ and signature $sig$ on WER and NSMD. The $c$ in the subscript stands for correct and $w$ stands for wrong. Only $f_{wm}$ with correct trigger and $sig$ can pass through verification.
@@
For $f_{wm}$, (i) when the trigger is wrong ($t_{w}$) and the signature is correct ($sig_{c}$), WER = 0, NSMD $>T_{N}$. This means that $f_{wm}$ has learned the relationship between $sig$ and $t$. Whether $t$ is correct determines whether $f_{wm}$ can produce the expected output, which in turn affects both the calculation of WER and NSMD. (ii) When the trigger is correct ($t_{c}$) and the signature is wrong (considering the most dangerous scenario that $sig_{w}$ consists only of $\{-1,1\}$), WER $\approx 0.5$. This is because $t_{c}$ leads to the right $sig^{\prime}$, and its expectation of WER with a random string $\{-1,1\}$ is $0.5$. As the output matrix is correctly generated by $f_{wm}$ based on $t_{c}$, NSMD = 00 in this case. (iii) When both trigger and signature are wrong ($t_{w}$ and $sig_{w}$), WER = 00, NSMD$>T_{N}$, indicating that the watermark cannot be correctly verified without providing the correct key. (iv) For a model without embedded watermarks $f_{clean}$, watermarks cannot be extracted even if the correct key is provided.
Table 2: WER results of different extractor $E$. $f_{wm}$, $E_{c}$, and $E_{w}$ means watermarked model, correct $E$ and wrong $E$, respectively. Only the correct E can accurately extract the signature.
@@
Wrong extractor $E$. Since $E$ is not involved in NSMD calculation, we only analyze the impact of $E$ on WER. As shown in Table 2, when $E$ is wrong, the WER is close to 00, indicating that wrong $E$ is unable to extract the watermark.
Table 3: NSMD of different null space matrix $N$. $f_{wm}$, $N_{c}$, $N_{r}$ and $N_{s}$ means watermarked model, correct $N$, random $N$ and $N$ composed of small elements, respectively. Only when the correct $N$ is used can the NSMD be close to 00.
@@
Table 4: Impact of fine-tuning on watermark performance. $F_{wm}$ means fine-tuned whole watermarked model and $F_{clean}$ denotes fine-tuned model without watermark. (i) $F_{wm}$ has a slight loss in ACC compared to $F_{clean}$. (ii) Fine-tuning has little effect on the WER of $F_{wm}$. (3) Fine-tuning increases the NSMD, but it is still significantly different from the model without watermark.
@@
Table 5: LL-LFEA results on NSmark. $f_{wm}$ means watermarked model and $f_{LL-LFEA}$ denotes $f_{wm}$ attacked by LL-LFEA. Due to the invariance of the null space to linear transformations, NSMD can still effectively prove the IP of the model despite the failure of WER.
Metric BERT RoBERTa DeBERTa XLNet
$f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$ $f_{wm}$ $f_{LL-LFEA}$
WER 1.00 0.27 1.00 0.07 1.00 0.27 1.00 0.00
Wrong null space $N$. Since $N$ is not involved in calculating WER, we only analyze its impact on NSMD. As shown in Table 3, $N_{r}$ is a randomly generated matrix with the same dimension as $N$ and each element is distributed between $[0,1]$. It can be seen that NSMD is very large at this time. However, if the attacker knows the watermark algorithm and want to reduce NSMD, a $N_{s}$ with extremely small elements might be generated. In this case, NSMD might meet the verification requirements. This indicates that NSMD cannot be used independently to verify the watermark.
4.3.3 Fidelity
We hope that NSmark  does not affect the performance on the original tasks. Thus, we add a downstream network to $f_{wm}$, and fine-tune the whole model $F_{wm}$ with the downstream dataset. $F_{clean}$ without watermark is fine-tuned as baseline. Table 4 shows that the watermark has almost no impact on the performance of the model on the original task.
@@ 4.3.4 Defense against LL-LFEA
Defense against LL-LFEA. When designing NSmark, we focus on resisting LL-LFEA and propose null space verification using NSMD. Table 5 shows the impact of LL-LFEA on watermark verification, where $f_{LL-LFEA}$ denotes the model $f_{wm}$ attacked by LL-LFEA. Experiments show that after LL-LFEA, WER drops significantly, as discussed in Section 3.1, but NSMD is still close to 00, verifying that NSMD is an effective indicator for LL-LFEA. Furthermore, after applying LL-LFEA, the attacker can add a network to $f_{LL-LFEA}$ and fine-tune it for downstream tasks (detailed results are presented in Appendix B.3). Additionally, since LL-LFEA causes minimal degradation in model performance, the attacker may attempt to further compromise the watermark through multiple LL-LFEA attacks. Analysis for this aspect is provided in Appendix B.7.
Recovery of WER. In LFEA (Li et al., 2023a), a method is proposed to recover the watermark. We revise this method to recover $f_{rec}$ from $f_{LL-LFEA}$. Specifically, assume that the output matrix of $f_{wm}$ is $A_{1(n\times m)}$ as Proof A.2. After being attacked with $Q_{(n\times n)}$, the output matrix turns to $A_{2}=Q\times A_{1}$. Therefore, an estimate of $Q$ can be obtained as $Q^{\prime}=A_{2}\times A_{1}^{-1}$. If $m\neq n$, then $A_{1}$ is not reversible and $Q^{\prime}=A_{2}\times A_{1}^{T}\times(A_{1}\times A_{1}^{T})^{-1}$. Then we perform an anti-attack transformation on $f_{LL-LFEA}$, that is, multiply all the outputs of $f_{LL-LFEA}$ by $Q^{\prime}$ to get $f_{rec}$. Figure 6 shows that after recovery, WER is significantly improved, indicating that such linear attacks are recoverable. In all cases, the NSMD is quite small, proving that NSMD is invariant to LL-LFEA. In the recovery algorithm in (Li et al., 2023a), both the attacker and the model owner might use such an algorithm to claim to be the owner of the model, which will cause verification ambiguity. However, our proposed NSMD is invariant under LL-LFEA, so as long as the timestamp information is added to the $key$ tuple, the ownership can be reliably verified according to the time sequence of the model and the release of $key$.
Figure 6: Changes of WER and NSMD before and after LL-LFEA attack and recovery. $f_{wm}$, $f_{LL-LFEA}$ and $f_{rec}$ denote watermarked model, $f_{wm}$ attacked by LL-LFEA, recovered model from $f_{LL-LFEA}$, respectively.
@@ 4.3.5 Robustness
The robustness of watermark refers to whether watermark can be effectively verified after watermark removal attacks. Next, we will analyze the robustness of NSmark  against fine-tuning, pruning, fine-pruning, and overwriting attacks. More robustness analysis against paraphrasing attack and multi-time LL-LFEA attack are shown in Appendix B.6B.7.
Robustness against fine-tuning. Table 4 shows the WER and NSMD results after fine-tuning on downstream tasks. $F_{wm}$ and $F_{clean}$ are obtained the same as in Section 4.3.3. In most cases, the WER is still very high, indicating that the embedded $sig$ can still be effectively extracted after downstream fine-tuning. However, the WER of RoBERTa on Lingspam task is relatively low. In main results we set the max input length to 128, which is quite shorter than the average length of Lingspam samples (average length of 695.26). Thus it is not sure the model’s input includes triggers (possibly truncated). We perform further experiments on increasing the max length of input to 512, and modify the position of trigger to the front. WER increases to more than $0.84$ and $0.87$ respectively. Besides, compared to $F_{clean}$, there is still obvious discrimination. Therefore, for complex tasks, the verification threshold $T_{W}$ can be slightly lowered.
Figure 7: Impact of pruning attacks on watermark. The dotted line is the performance of the original model without watermark.
Robustness against pruning and fine-pruning. Pruning is a commonly used model compression method and is often used to destroy the watermark embedded in the model. Referring to (Han et al., 2015; Shao et al., 2024), we sort the parameters of each layer in LM, then set some fractions of parameters with smaller absolute value to 00. Figure 7 shows that when the pruning rate is less than or equal to $0.8$, the WER is close to $1.0$. When the pruning rate is less than or equal to $0.6$, NSMD does not change significantly, and even when the pruning rate is as high as $0.9$, NSMD is still distinguishable. Besides, as shown in Appendix B.4, the accuracy of the watermarked model only changes slightly after pruning then fine-tuning on the SST-5. This shows that the embedded watermark is robust to pruning attack.
Usually, pruning will affect the performance of the model on the original task, and the original task accuracy will be restored by fine-tuning (fine-pruning), as demonstrated by the results in Appendix B.5.
Robustness against overwriting. Overwriting means attacker embeds his own watermark into a model that has already been watermarked in the same way. This may destroy the original watermark. We simulate this process to obtain $f_{ow}$, then add a downstream network and fine-tune to obtain $F_{ow}$. We test the original watermark as shown in Table 6. The overwriting attack has little effect on ACC and WER except on Lingspam. Meanwhile, it has an impact on NSMD similar to that of fine-tuning.
Table 6: Impact of overwriting attacks on watermark performance. $f_{ow}$ is the overwritten model, which is fine-tuned to obtain $F_{ow}$. “–” means not applicable.